Consider the function y=\ ln(1+x).

Which of the following is the Maclaurin series expansion of the first three terms?


  1. "x-\frac{x^2}{2}+\frac{x^3}{6}"
  2. "x-\frac{x^2}{3}+\frac{x^3}{2}"
  3. "x-\\frac{x^2}{2}+\\frac{x^3}{3}\u200b"
  4. "x+\frac{x^2}{2}-\frac{x^3}{3}"
1
Expert's answer
2021-09-23T16:54:25-0400

We should obtain the first three terms of the series "{\displaystyle f(x)=\sum _{n=0}^{+\infty }{\frac {f^{(n)}(0)}{n!}}x^{n}}"

"f^{(0)}(0) = \ln (1+0) = 0, \\
f^{(1)}(0) = \frac{1}{1+0} = 1, \\
f^{(2)}(0) = -\frac{1}{(1+0)^2} = -1, \\
f^{(3)}(0) = \frac{2}{(1+0)^3} = 2"


"\ln(1+x) = 0 + x - \frac{1}{2}x^2 + \frac{2}{6}x^3 = x - \frac{1}{2}x^2 + \frac{1}{3}x^3." The correct answer is 3